# Synthetic Object Compositions for Scalable and Accurate Learning in Detection, Segmentation, and Grounding

### v3

@@ 1 Introduction
Human perception is inherently structured: we naturally organize the visual world into coherent regions and objects based on principles of proximity, similarity, continuity, and closure—phenomena first formalized in Gestalt psychology [wertheimer1923untersuchungen, kanizsa1979organization]. In computer vision, tasks such as object detection, instance segmentation, and referring expression grounding operationalize this capability and power a broad range of applications, from robotic perception and autonomous driving to photo editing and assistive technologies.
@@
Historically, the success of segmentation and detection models remains heavily dependent on the quality, scale, and diversity of annotated datasets [Lin2014COCO, Cordts2016Cityscapes, Gupta_2019_CVPR]. Existing annotated datasets are painstaking and expensive to expand. For example, COCO [lin2015microsoftcococommonobjects] required 2.2M worker hours to annotate only 100K images with 80 object categories. They are also biased toward a limited number of categories, and often fail to cover the full combinatorial diversity of scenes, objects, and referring expressions encountered in the real world.
@@
To overcome these challenges, researchers have turned to synthetic data. Synthetic datasets by rendering scenes in simulation [Jiang2022PseudoQGP, Peng2023Kosmos2GM, Wang2024LearningVG, he2024learningsyntheticdatavisual, zhang2025scaling, kirillov2023segment, wu2023datasetdm, yang2023freemask, ye2024seggen, zhang2024maskfactory, zhao2023xpaste, masi2024harlequin, Wrenninge2018Synscapes, Cabon2020VKITTI2, Roberts_2021_ICCV, Liu2019CLEVRRef] allow for full control and offer accurate dense ground truth at a large scale, but are still geared toward rigid domains like indoor or driving scenes and are limited in object diversity due to the scarcity of 3D object assets. On the other hand, model-annotated synthetic datasets on real or generated images [Jiang2022PseudoQGP, Peng2023Kosmos2GM, Wang2024LearningVG, he2024learningsyntheticdatavisual, zhang2025scaling, kirillov2023segment, wu2023datasetdm, yang2023freemask, ye2024seggen, zhang2024maskfactory, zhao2023xpaste] introduce greater variability in scene composition and object appearance, yet they inherit annotation noise from both the labeling models and the image generators.
@@
In this work, we introduce SOC (Figure 1): a scalable data synthesis pipeline that departs from prior paradigms. Rather than rendering entire scenes or relying on models to annotate images, we employ an object-centric composition strategy: we first collect high-quality synthetic object segments using a strong generative model; next, we compose new images by pasting segments according to designed layouts with 3D geometric layout augmentation and camera configuration augmentation to ensure diversity and robustness; finally, we apply generative harmonization with mask-area-weighted blending leveraging recent advances [zhang2025scaling] to enhance the images’ photorealism. SOC not only provides controllability and accurate region annotations like simulation-based methods, but also offers the composition flexibility and scene diversity of model-annotated approaches. Specifically, we generate 20M synthetic object segments; using these, one can compose any number of synthetic images, each with corresponding mask, box, label, and referring expression annotations.
@@
In open‐vocabulary detection and visual grounding, we show that SOC is an efficient and scalable approach to improve model performance, and complementary to real datasets. Even a small SOC dataset of 50K images delivers strong gains: it lifts LVIS [gupta2019lvisdatasetlargevocabulary] AP from 20.1 to 29.8 (+9.7) and rare‐class AP from 10.1 to 23.5 (+13.4), exceeding the +7.0 boost from 20M model generated data (GRIT [Wu2022GRiTAG]) and matching the +10.5 gain from 200K human-annotated data (V3Det [zhao2024open, Wu2022GRiTAG, wang2023v3detvastvocabularyvisual]). Scaling SOC from 50K to 400K further boosts LVIS AP to 31.4 (+1.6) and OdinW‐35 mAP to 22.8 (+1.8), demonstrating the scalability of SOC. Additionally, SOC complements existing real datasets with more balanced category distribution, wider coverages, and compositions: adding 100K SOC images above model-generated data (GRIT) and human-annotated data (V3Det) still yields additive gains of +6.2 rare‐class LVIS AP, and +2.8 average AP on OdinW‐35.
@@ 2 Related work
Datasets for object detection, segmentation, and referring expression. Detection and segmentation benchmarks include COCO [lin2015microsoftcococommonobjects], OpenImages [Kuznetsova_2020], Object365 [9009553], LVIS [gupta2019lvisdatasetlargevocabulary], V3Det [wang2023v3detvastvocabularyvisual], and ODInW [li2021grounded, li2022elevater]. Semantic segmentation datasets include COCO‐Stuff [Caesar2016COCOStuffTA], ADE20K [8100027, zhou2018semanticunderstandingscenesade20k], PASCAL VOC [everingham2011pascal], PASCAL Context [Mottaghi2014TheRO], and Cityscapes [Cordts2016Cityscapes]. Referring expression benchmarks include ReferItGame [kazemzadeh-etal-2014-referitgame], RefCOCO/+/g [Yu2016ModelingCI, Mao2015GenerationAC], Flickr30K Entities [plummer2016flickr30kentitiescollectingregiontophrase], Visual Genome [Krishna2016VisualGC], and GoldG [li2022groundedlanguageimagepretraining].
Synthetic data generation pipelines. Copy-paste methods [Dvornik2018ModelingVC, Fang2019InstaBoostBI, Ghiasi2020SimpleCI] and X-Paste [Zhao2022XPasteRS] paste segments onto backgrounds but lack photorealism. Diffusion-based approaches including ControlNet [Zhang2023AddingCC], LayoutDiffusion [Zheng2023LayoutDiffusionCD], MaskFactory [zhang2024maskfactory], SegGen [ye2024seggen], and DatasetDM [wu2023datasetdm] generate images from masks or layouts but often produce inaccurate annotations. Pseudo-labeling methods including GRIT [gupta2022gritgeneralrobustimage], SynGround [he2024learningsyntheticdatavisual], and Learning VG [Wang2024LearningVG] apply detectors to real or generated images. In contrast, SOC uses object composition with 3D geometric layout augmentation and generative harmonization for pixel-perfect annotations.
Image matting, blending, and harmonization. Image matting methods including Deep Image Matting [Xu2017DeepIM], attention-guided matting [Qiao2020AttentionGuidedHS], Human Instance Matting [Sun2022HumanIM], Referring Image Matting [Li2023ReferringIM], DFIMat [Jiao2024DFIMatDF], MP-Mat [Jiao2025MPMatA3H], and DIS [qin2022highlyaccuratedichotomousimage] extract object segments with precise alpha mattes. We use DIS for segment extraction and IC-Light [zhang2025scaling] for diffusion-based harmonization, combined with mask-area-weighted blending for photorealism.
@@ 3.1 Synthesizing object segments
Given 46000+ object categories we collected, we and prompt Qwen 2.5-32B [qwen2025qwen25technicalreport] to produce text descriptions for each of them. These prompts are then passed to FLUX‑1‑dev [flux2024], a strong text-to-image generation model, yielding single‑object images rendered on a uniform white background with randomly sampled viewpoints to achieve diverse visual info for each objects.. Finally, we apply DIS [qin2022highlyaccuratedichotomousimage] to extract object segments from white-background images. We find that this setup produces cleaner mask boundaries than generating objects in cluttered scenes.
@@
Models trained on real datasets often exploit spurious correlations between object categories and pictorial cues (e.g., “cars appear large and near the bottom”) [chen2020monopair], rather than learning robust semantic features. To break these shortcuts, we adopt a 3D geometric layout augmentation strategy: we model each composite image as a 3D scene where depth and spatial position are sampled independently of object category, i.e., $p(d_{i},X_{i},Y_{i}|c_{i})=p(d_{i},X_{i},Y_{i})$. This ensures objects of the same category appear at diverse depths, sizes, and positions, preventing category-specific pictorial patterns. For each image, we sample 5-20 object segments (matching COCO/SA-1B distributions) using balanced category sampling to avoid bias.
@@ 3.3 Generative harmonization
Empirically, we find that naively pasting objects onto a background image often produces unnaturally sharp edges; segmentation models trained on these data can exploit these edge artifacts rather than truly learning semantics. To mitigate this shortcut, we employ IC‑Light [zhang2025scaling], a diffusion‑based model that simultaneously performs background inpainting and global relighting to harmonize the composition. IC-Light generates a compatible background for the pasted objects and adjusts lighting conditions across the entire scene, producing more photorealistic images without simple edge‑based artifacts. However, IC-Light sometimes distorts fine details of smaller objects or alters object colors (e.g. blue $\rightarrow$ red), breaking consistency with text descriptions. To address this, we re-blend the original segments with the harmonized images using a blending weight $\alpha_{i}\in[0,1]$ for each object mask $M_{i}$, where smaller objects receive higher $\alpha_{i}$ to preserve their original appearance. Finally, we apply a lightweight soft matting step that converts the binary masks into soft alpha mattes to smooth object boundaries and better integrate the foregrounds into the harmonized background. This blending procedure yields a 2.3% AP boost on LVIS-mini-val (Fig. 3).
@@
Human-annotated COCO [lin2015microsoftcococommonobjects], LVIS [gupta2019lvisdatasetlargevocabulary], V3Det [wang2023v3detvastvocabularyvisual] Limited $\sim$1M Sec 4.14.5
Real Image + Pseudo annotator GRIT [gupta2022gritgeneralrobustimage] ✗ (bbox only) $\sim$20M Sec 4.1, 4.2, 4.5
Diffusion Image + Pseudo annotator SynGround [he2024learningsyntheticdatavisual] ✗ (pseudo) Limited Unlimited (current 1.2M) Sec 4.6
Mask/Layout control Diffusion Image
MaskFactory [zhang2024maskfactory], SegGen [ye2024seggen]
Limited Limited Limited Unlimited (current 1M) Sec 4.6
Simple-Copy-Paste Simple-Copy-Paste [Ghiasi2020SimpleCI], X-Paste [Zhao2022XPasteRS] Limited ($\sim$1K) Depends on real background Sec 4.6
@@ 3.5 Generating region annotations
We generate annotations for object detection (OD), instance segmentation (SEG), and visual grounding (VG). For OD and SEG, we compute bounding boxes and masks by subtracting occluded pixels from each object’s original mask. For VG, we prompt QwQ-32B [qwq32b] with each object’s bounding box, mask, category, and generation prompt to produce 3-6 attribute-based and spatial-based referring expressions per type, yielding at least 9 dense expressions per image.
@@ 4.1 Task 1: Open-vocabulary object detection
Adding various scale SOS datasets.
@@
Combining human (V3Det) and model (GRIT) annotated datasets with SOS datasets.
@@
Data Data Scale gRefCOCO DoD (FULL) RefCOCO avg. P@1 (F1=1, IoU0.50.5) Nacc mAP mAPshort mAPmid mAPlong mAPvery_long P@1 P@5 P@10 O365+GoldG (Baseline) 1.4M 39.8 89.3 15.6 17.3 16.7 14.3 13.1 54.3 89.3 94.6 Adding human (GRIT) or model (V3Det) annotated datasets. +GRIT +20M 40.7 89.3 17.0 17.7 18.0 15.7 15.7 56.5 90.5 95.4 +V3Det +200K 40.3 89.3 16.7 17.7 18.0 15.7 15.7 55.0 89.4 94.6 Adding various‐scale SOS datasets. +SOC-FC-50K +50K 41.2 93.9 16.6 18.3 17.1 15.7 14.5 54.3 90.3 95.1 +SOC-FC-100K +100K 41.3 97.7 19.4 22.0 20.4 17.8 16.9 54.0 90.4 94.9 +SOC-FC-50K+SOC-GC-50K +100K 40.9 93.0 17.3 20.0 18.1 15.7 14.8 55.0 89.9 94.9 Combining human, model‐annotated, and SOS datasets. O365+GoldG+GRIT+V3Det 21.6M 41.0 89.3 17.5 23.4 18.3 14.7 13.8 56.4 90.9 95.8 +GRIT+V3Det+SOC-FC-50K+SOC-GC-50K +100K 40.9 94.4 18.9 23.7 20.5 15.9 14.3 55.1 89.8 94.8
To study the efficacy of SOC data on open-vocabulary object detection, we use MM-Grounding-DINO [zhao2024open], a strong open‐vocabulary detector and visual grounding model that handles various types of textual input and provides bounding boxes. We initialize the model with real-data pretrained weights and continue training on our synthetic dataset. We report performance on two benchmarks: (1). LVIS [gupta2019lvisdatasetlargevocabulary] for open‐vocabulary detection, reporting AP on the LVIS 1.0 Full Val split (including APrare, APcommon, and APfrequent); (2). OdinW-35 [li2021grounded, li2022elevater] for scene‐specific detection, evaluating across 35 diverse scenarios and reporting the average AP.
Baselines. We compare with 4 baseline training datasets: (1). Object365+GoldG: trains on Object365 [9009553] (600K images, 365 open‐vocabulary categories) together with GoldG [li2022groundedlanguageimagepretraining] (800 K vision–grounding examples mixing GQA [hudson2018gqa] and Flickr30K Entities [plummer2016flickr30kentitiescollectingregiontophrase]); (2). Object365+GoldG+GRIT: adds GRIT [schuhmann2022laion5bopenlargescaledataset], a 20M example visual grounding dataset curated from LAION-2B [ziyiLaion-2B] and COYO-700M [kakaobrain2022coyo-700m]; (3). Object365+GoldG+V3Det: augments Object365+GoldG with V3Det [wang2023v3detvastvocabularyvisual], a 200k human‐annotated detection dataset covering 13K categories; and (4). Object365+GoldG+V3Det+GRIT: combines all four real datasets [gupta2022gritgeneralrobustimage].
@@ 4.2 Task 2: Visual grounding
Figure 4: (Sec. 4.3) SOS consistently improves LVIS instance segmentation, with the largest gains on rare categories.
@@ 4.4 Task 4: Small‐vocabulary, limited‐data regimes
Real-world segmentation applications (e.g., surveillance cameras) only need a small set of categories and have limited annotation budgets for the dataset curation, in this section, we evaluate SOC specifically in a small-vocabulary, limited-data regime to investigate its effectiveness under these practical constraints. We adopt Mask2Former [cheng2022maskedattentionmasktransformeruniversal], a widely-used closed-vocabulary segmentation model, for the instance segmentation task. Specifically, we compare a Mask2Former-ResNet-50 [he2016deep] baseline trained solely on COCO instance-segmentation subsets across $4$ different data scales (1K 1%), (10K (10%), 50K (50%), or all COCO images) against a setting augmented with our SOC data. This augmented data is generated by mixing real segments collected from corresponding splits and synthetic segments from SOC synthetic segments within 80 COCO categories.

### v4

@@ 1 Introduction
Human perception is inherently structured: we naturally organize the visual world into coherent regions and objects based on principles of proximity, similarity, continuity, and closure—phenomena first formalized in Gestalt psychology [92, 41]. In computer vision, tasks such as object detection, instance segmentation, and referring expression grounding operationalize this capability and power a broad range of applications, from robotic perception and autonomous driving to photo editing and assistive technologies.
@@
Historically, the success of segmentation and detection models remains heavily dependent on the quality, scale, and diversity of annotated datasets [60, 10, 24]. Existing annotated datasets are painstaking and expensive to expand. For example, COCO [59] required 2.2M worker hours to annotate only 100K images with 80 object categories. They are also biased toward a limited number of categories, and often fail to cover the full combinatorial diversity of scenes, objects, and referring expressions encountered in the real world.
@@
To overcome these challenges, researchers have turned to synthetic data. Synthetic datasets by rendering scenes in simulation [37, 70, 91, 30, 111, 43, 95, 98, 102, 72, 93, 3, 80, 62] allow for full control and offer accurate dense ground truth at a large scale, but are still geared toward rigid domains like indoor or driving scenes and are limited in object diversity due to the scarcity of 3D object assets. On the other hand, model-annotated synthetic datasets on real or generated images [37, 70, 91, 30, 111, 43, 95, 98, 102, 72] introduce greater variability in scene composition and object appearance, yet they inherit annotation noise from both the labeling models and the image generators.
@@
In this work, we introduce SOC (Figure 1): a scalable data synthesis pipeline that departs from prior paradigms. Rather than rendering entire scenes or relying on models to annotate images, we employ an object-centric composition strategy: we first collect high-quality synthetic object segments using a strong generative model; next, we compose new images by pasting segments according to designed layouts with 3D geometric layout augmentation and camera configuration augmentation to ensure diversity and robustness; finally, we apply generative harmonization with mask-area-weighted blending leveraging recent advances [111] to enhance the images’ photorealism. SOC not only provides controllability and accurate region annotations like simulation-based methods, but also offers the composition flexibility and scene diversity of model-annotated approaches. Specifically, we generate 20M synthetic object segments; using these, one can compose any number of synthetic images, each with corresponding mask, box, label, and referring expression annotations.
@@
In open‐vocabulary detection and visual grounding, we show that SOC is an efficient and scalable approach to improve model performance, and complementary to real datasets. Even a small SOC dataset of 50K images delivers strong gains: it lifts LVIS [25] AP from 20.1 to 29.8 (+9.7) and rare‐class AP from 10.1 to 23.5 (+13.4), exceeding the +7.0 boost from 20M model generated data (GRIT [94]) and matching the +10.5 gain from 200K human-annotated data (V3Det [113, 94, 90]). Scaling SOC from 50K to 400K further boosts LVIS AP to 31.4 (+1.6) and OdinW‐35 mAP to 22.8 (+1.8), demonstrating the scalability of SOC. Additionally, SOC complements existing real datasets with more balanced category distribution, wider coverages, and compositions: adding 100K SOC images above model-generated data (GRIT) and human-annotated data (V3Det) still yields additive gains of +6.2 rare‐class LVIS AP, and +2.8 average AP on OdinW‐35.
@@ 2 Related work
Datasets for object detection, segmentation, and referring expression. Detection and segmentation benchmarks include COCO [59], OpenImages [47], Object365 [84], LVIS [25], V3Det [90], and ODInW [56, 51]. Semantic segmentation datasets include COCO‐Stuff [4], ADE20K [118, 119], PASCAL VOC [16], PASCAL Context [68], and Cityscapes [10]. Referring expression benchmarks include ReferItGame [42], RefCOCO/+/g [105, 66], Flickr30K Entities [71], Visual Genome [46], and GoldG [54].
Synthetic data generation pipelines. Copy-paste methods [15, 18, 20] and X-Paste [112] paste segments onto backgrounds but lack photorealism. Diffusion-based approaches including ControlNet [110], LayoutDiffusion [116], MaskFactory [72], SegGen [102], and DatasetDM [95] generate images from masks or layouts but often produce inaccurate annotations. Pseudo-labeling methods including GRIT [26], SynGround [30], and Learning VG [91] apply detectors to real or generated images. In contrast, SOC uses object composition with 3D geometric layout augmentation and generative harmonization for pixel-perfect annotations.
Image matting, blending, and harmonization. Image matting methods including Deep Image Matting [97], attention-guided matting [73], Human Instance Matting [87], Referring Image Matting [52], DFIMat [38], MP-Mat [39], and DIS [74] extract object segments with precise alpha mattes. We use DIS for segment extraction and IC-Light [111] for diffusion-based harmonization, combined with mask-area-weighted blending for photorealism.
@@ 3.1 Synthesizing object segments
Given 46000+ object categories we collected, we and prompt Qwen 2.5-32B [75] to produce text descriptions for each of them. These prompts are then passed to FLUX‑1‑dev [48], a strong text-to-image generation model, yielding single‑object images rendered on a uniform white background with randomly sampled viewpoints to achieve diverse visual info for each objects.. Finally, we apply DIS [74] to extract object segments from white-background images. We find that this setup produces cleaner mask boundaries than generating objects in cluttered scenes.
@@
Models trained on real datasets often exploit spurious correlations between object categories and pictorial cues (e.g., “cars appear large and near the bottom”) [7], rather than learning robust semantic features. To break these shortcuts, we adopt a 3D geometric layout augmentation strategy: we model each composite image as a 3D scene where depth and spatial position are sampled independently of object category, i.e., $p(d_{i},X_{i},Y_{i}|c_{i})=p(d_{i},X_{i},Y_{i})$. This ensures objects of the same category appear at diverse depths, sizes, and positions, preventing category-specific pictorial patterns. For each image, we sample 5-20 object segments (matching COCO/SA-1B distributions) using balanced category sampling to avoid bias.
@@ 3.3 Generative harmonization
Empirically, we find that naively pasting objects onto a background image often produces unnaturally sharp edges; segmentation models trained on these data can exploit these edge artifacts rather than truly learning semantics. To mitigate this shortcut, we employ IC‑Light [111], a diffusion‑based model that simultaneously performs background inpainting and global relighting to harmonize the composition. IC-Light generates a compatible background for the pasted objects and adjusts lighting conditions across the entire scene, producing more photorealistic images without simple edge‑based artifacts. However, IC-Light sometimes distorts fine details of smaller objects or alters object colors (e.g. blue $\rightarrow$ red), breaking consistency with text descriptions. To address this, we re-blend the original segments with the harmonized images using a blending weight $\alpha_{i}\in[0,1]$ for each object mask $M_{i}$, where smaller objects receive higher $\alpha_{i}$ to preserve their original appearance. Finally, we apply a lightweight soft matting step that converts the binary masks into soft alpha mattes to smooth object boundaries and better integrate the foregrounds into the harmonized background. This blending procedure yields a 2.3% AP boost on LVIS-mini-val (Fig. 3).
@@
Human-annotated COCO [59], LVIS [25], V3Det [90] Limited $\sim$1M Sec 4.14.5
Real Image + Pseudo annotator GRIT [26] ✗ (bbox only) $\sim$20M Sec 4.1, 4.2, 4.5
Diffusion Image + Pseudo annotator SynGround [30] ✗ (pseudo) Limited Unlimited (current 1.2M) Sec 4.6
Mask/Layout control Diffusion Image MaskFactory [72], SegGen [102] Limited Limited Limited Unlimited (current 1M) Sec 4.6
Simple-Copy-Paste Simple-Copy-Paste [20], X-Paste [112] Limited ($\sim$1K) Depends on real background Sec 4.6
@@ 3.5 Generating region annotations
We generate annotations for object detection (OD), instance segmentation (SEG), and visual grounding (VG). For OD and SEG, we compute bounding boxes and masks by subtracting occluded pixels from each object’s original mask. For VG, we prompt QwQ-32B [89] with each object’s bounding box, mask, category, and generation prompt to produce 3-6 attribute-based and spatial-based referring expressions per type, yielding at least 9 dense expressions per image.
@@ 4.1 Task 1: Open-vocabulary object detection
Adding various scale SOC datasets.
@@
Combining human (V3Det) and model (GRIT) annotated datasets with SOC datasets.
@@
Data Data Scale gRefCOCO DoD (FULL) RefCOCO avg. P@1 (F1=1, IoU0.50.5) Nacc mAP mAPshort mAPmid mAPlong mAPvery_long P@1 P@5 P@10 O365+GoldG (Baseline) 1.4M 39.8 89.3 15.6 17.3 16.7 14.3 13.1 54.3 89.3 94.6 Adding human (GRIT) or model (V3Det) annotated datasets. +GRIT +20M 40.7 89.3 17.0 17.7 18.0 15.7 15.7 56.5 90.5 95.4 +V3Det +200K 40.3 89.3 16.7 17.7 18.0 15.7 15.7 55.0 89.4 94.6 Adding various‐scale SOC datasets. +SOC-FC-50K +50K 41.2 93.9 16.6 18.3 17.1 15.7 14.5 54.3 90.3 95.1 +SOC-FC-100K +100K 41.3 97.7 19.4 22.0 20.4 17.8 16.9 54.0 90.4 94.9 +SOC-FC-50K+SOC-GC-50K +100K 40.9 93.0 17.3 20.0 18.1 15.7 14.8 55.0 89.9 94.9 Combining human, model‐annotated, and SOC datasets. O365+GoldG+GRIT+V3Det 21.6M 41.0 89.3 17.5 23.4 18.3 14.7 13.8 56.4 90.9 95.8 +GRIT+V3Det+SOC-FC-50K+SOC-GC-50K +100K 40.9 94.4 18.9 23.7 20.5 15.9 14.3 55.1 89.8 94.8
To study the efficacy of SOC data on open-vocabulary object detection, we use MM-Grounding-DINO [113], a strong open‐vocabulary detector and visual grounding model that handles various types of textual input and provides bounding boxes. We initialize the model with real-data pretrained weights and continue training on our synthetic dataset. We report performance on two benchmarks: (1). LVIS [25] for open‐vocabulary detection, reporting AP on the LVIS 1.0 Full Val split (including APrare, APcommon, and APfrequent); (2). OdinW-35 [56, 51] for scene‐specific detection, evaluating across 35 diverse scenarios and reporting the average AP.
Baselines. We compare with 4 baseline training datasets: (1). Object365+GoldG: trains on Object365 [84] (600K images, 365 open‐vocabulary categories) together with GoldG [54] (800 K vision–grounding examples mixing GQA [35] and Flickr30K Entities [71]); (2). Object365+GoldG+GRIT: adds GRIT [83], a 20M example visual grounding dataset curated from LAION-2B [123] and COYO-700M [2]; (3). Object365+GoldG+V3Det: augments Object365+GoldG with V3Det [90], a 200k human‐annotated detection dataset covering 13K categories; and (4). Object365+GoldG+V3Det+GRIT: combines all four real datasets [26].
@@ 4.2 Task 2: Visual grounding
Figure 4: (Sec. 4.3) SOC consistently improves LVIS instance segmentation, with the largest gains on rare categories.
@@ 4.4 Task 4: Small‐vocabulary, limited‐data regimes
Real-world segmentation applications (e.g., surveillance cameras) only need a small set of categories and have limited annotation budgets for the dataset curation, in this section, we evaluate SOC specifically in a small-vocabulary, limited-data regime to investigate its effectiveness under these practical constraints. We adopt Mask2Former [8], a widely-used closed-vocabulary segmentation model, for the instance segmentation task. Specifically, we compare a Mask2Former-ResNet-50 [28] baseline trained solely on COCO instance-segmentation subsets across $4$ different data scales (1K 1%), (10K (10%), 50K (50%), or all COCO images) against a setting augmented with our SOC data. This augmented data is generated by mixing real segments collected from corresponding splits and synthetic segments from SOC synthetic segments within 80 COCO categories.